40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 28 04:27:33 2016, ericq, Update, General, mucking about alscheck.pdfweird_olpit.pdf
    Reply  Thu Sep 29 09:49:53 2016, rana, Update, General, mucking about 
    Reply  Sat Oct 1 10:03:28 2016, ericq, Update, General, mucking about 
       Reply  Tue Oct 4 02:59:48 2016, ericq, Update, General, mucking about 2016-10-04_ALSXspectra.pdf
Message ID: 12521     Entry time: Wed Sep 28 04:27:33 2016     Reply to this: 12522   12527
Author: ericq 
Type: Update 
Category: General 
Subject: mucking about 

PMC was terribly misaligned. The PMCR camera seems to have drifted somewhat off target too, but I didn't touch it.

Realigned ITMX for the nth time today.

Finding ALSY beatnote was easy, ALSX eludes me. I did a rough one-point realignment on the X beat PD which is usually enough, but it's probably been long enough that near/far field alignmnet is neccesary. 

ALSY noise is mostly nominal, but there is a large 3Hz peak that is visible in the spot motion, and also modulates the beat amplitude by multiple dBs.

It looked to me that the ETMY oplev spot was moving too much, which led me to measure the oplev OLGs. There is some wierd inter-loop interference going on between OLPIT and OLYAW. With both on (whether OSEM damping is on or off, so input matrix shenanigans can't be to blame) there is a very shallow "notch" at around 4.5Hz, which leads to very little phase at 3Hz, and thus tons of control noise. Turning the OL loop not being measured off makes this dip go away, but the overall phase is still signfinicantly less than we should have. I'm not sure why. I'll just show the PIT plot, but things look pretty much the same for YAW. 


I did some more ETMX tests. Locked arm, raised the servo output limit to 15k, then increased the gain to make the loop unstable. I saw the SUS LSC signals go up to tens of thousands of counts when the unlock happened. I did this a dozen times or so, and every time the ETM settled in the same angular position according to the oplev.

Right now, another hysteresis script is running, misaliging in pitch and yaw. Amplitude 1V in each direction. So far, everything is stable after three on/off cycles.

Attachment 1: alscheck.pdf  103 kB  | Hide | Hide all
alscheck.pdf
Attachment 2: weird_olpit.pdf  5 kB  | Hide | Hide all
weird_olpit.pdf
ELOG V3.1.3-